2HI -> H2 I2

Keq= 49.5 at 440'C


If 5 0. M HI is initially placed into a container, what will be the equilibrium [ ] HI ?


1
Expert's answer
2020-01-24T05:00:09-0500

"K_{eq}=\frac{[H_2][I_2]}{[HI]^2}"

"2HI \to H_2+I_2\\"

"5-x" "\frac{x}{2}" "\frac{x}{2}"

The initian concentration is "5 M."

"K_{eq}=49.5" "=\frac{\frac{x}{2}\times\frac{x}{2}}{(5-x)^2}"

"\implies197x^2-1980x+4950=0"

"\implies x=4.68"

The equilibrium concentration of "[HI]" is "=5-x=5-4.68=0.32\ M"